# Assessing the Limitations of Large Language Models in Clinical Fact Decomposition

### v1

@@ 3.1 RQ1 — Fact Decomposition
Figure 4: Distribution of coefficient of variation across documents. Higher values indicate higher variance in the number of facts per documents across LLMs.
@@
Using Earth Mover’s Distance (EMD) with a cosine cost function, we found that GPT-4o and o1-mini produced the most similar fact decompositions, closely followed by Gemini-1.5, whose pairwise EMD values against these models were consistently within 0.02 units of each other (Figure 3). In contrast, Llama3’s decompositions were notably more divergent, with pairwise EMD scores roughly 0.03–0.05 units higher. Compared to the original source sentences, GPT-4o decompositions had the lowest mean EMD, while Llama3’s were highest. Differences were also influenced by note type: procedure notes showed minimal variation due to their structured format, nursing notes had slightly higher EMDs reflecting more dynamic content, and discharge summaries—being lengthy and complex—exhibited the highest EMD values.
Figure 5: Sentence atomicity by note type and LLM, shown as the normalized distribution of the estimated number of facts-per-sentence.
@@
Sentence Atomicity
Overall, o1-mini produced the most facts per sentence, followed closely by GPT-4o, both maintaining similar rates of zero-fact sentences (Figure 5). Gemini-1.5 had more zero-fact sentences than GPT-4o and o1-mini, but fewer than Llama3. Llama3 produced facts for only about 20% of sentences in progress notes and discharge summaries, and, except for nursing notes, generated fewer facts than sentences (Table 1).
Qualitative Expert Evaluation
Expert reviewers assessed overall completeness of fact decompositions and the correctness, independence, and atomicity of individual facts. Completeness, defined as capturing all information in the source, is low: only 5% (1/20) of decompositions meet this criterion. For individual facts, reviewers judge whether each fact accurately reflected the source (correctness), stood on its own (independence), and represented a minimal unit of information (atomicity).
Across models, Llama3 had the highest rate of incorrect facts. Discharge summaries yield the least number of atomic facts, reflecting their complexity, while progress and procedure notes contain more granular and independent facts. Detailed results for GPT-4o are shown in Figure 6; additional results for other models and a summary of key metrics are reported in Appendix H and Table 2, respectively.
@@
Procedure Note Nursing Note
$P*$ $R*$ $R*_{w}$ $F1*$ $F1*_{w}$ $P*$ $R*$ $R*_{w}$ $F1*$ $F1*_{w}$
GPT-4o 98.5 78.7 92.4 86.5 95.0 97.5 88.9 92.2 92.3 94.2
o1-mini 97.8 78.4 93.0 86.2 95.4 96.6 88.5 93.3 91.6 94.4
Gemini-1.5 95.9 64.2 85.7 77.0 91.7 93.2 77.3 83.7 84.4 88.3
Llama3-8B 84.2 49.4 73.7 62.0 82.3 84.1 56.6 69.9 65.1 75.6
Progress Note Discharge Summary
$P*$ $R*$ $R*_{w}$ $F1*$ $F1*_{w}$ $P*$ $R*$ $R*_{w}$ $F1*$ $F1*_{w}$
GPT-4o 97.4 78.1 85.2 85.8 89.9 97.0 79.0 86.9 86.8 91.1
o1-mini 95.2 79.7 85.7 86.6 89.7 94.8 81.1 88.2 86.7 90.0
Gemini-1.5 96.0 66.2 75.2 77.9 83.9 96.5 65.4 76.3 76.9 83.8
Llama3-8B 83.4 41.9 57.4 53.8 67.0 84.6 38.9 57.3 50.4 67.2
[…]
3.2 RQ2 — Entailment
GPT-4o achieve the best F1 scores on four of the five benchmark datasets, including the FactEHR development set (full set of results in Appendix G, Table 6), and was therefore selected as the entailment judge for all 987,266 entailment pairs in the FactEHR dataset. From these, we randomly sampled 1,036 pairs for manual annotation by seven clinical experts. On doubly annotated entailment pairs, Fleiss’ Kappa was 0.67, indicating substantial inter-rater agreement. Under this evaluation, GPT-4o achieves a recall of 0.96 and precision of 0.86 (Appendix G, Table 7), closely mirroring its performance on the FactEHR development set. Overall, these findings suggest that GPT-4o serves as a reasonable proxy for human judgment in entailment assessments.
3.3 RQ3 — Fact Verification
Across all note types, fact precision remained consistently higher and less variable than fact recall—indicating that while the generated facts were generally accurate, many relevant facts were omitted, resulting in greater variability in recall and aligning with our qualitative findings (Appendix G, Figure 7).
4 Related Work
Because LLMs encode extensive clinical and biomedical knowledge (Singhal et al., 2023a; Agrawal et al., 2022; Munnangi et al., 2024), their evaluation is challenging. Models may not follow desired formats or may produce semantically equivalent but lexically divergent responses (Wadhwa et al., 2023; Agrawal et al., 2022). Prior work (Min et al., 2023; Xie et al., 2024; Tian et al., 2023) attempts to automatically assess factuality and completeness by decomposing text into atomic facts; for example, (Xie et al., 2024) use atomic facts for claim precision and recall calculations via entailment models. However, it remains unclear how effectively current models generate accurate, comprehensive atomic facts for clinical text.
While considerable progress has been made on atomic fact generation in general domains (Min et al., 2023) and biomedical literature (Wadden et al., 2020; Wright et al., 2022), there is little to no research focusing on clinical notes. Existing frameworks such as MAIRA-2 (Bannur et al., 2024) address fact-checking in radiology reports but not the broader range of clinical documents. Another, recent work has evaluated the role of context in fact decomposition, arguing for “molecular facts” that balance granularity with sufficient context. (Gunjal and Durrett, 2024).
5 Conclusion
@@
In this study, we evaluate four LLMs’ ability to decompose over 2,000 real-world, EHR-derived clinical notes of four types into independent facts. We examine the length and similarity of the decompositions and employed GPT-4o as an entailment judge to assess their accuracy and completeness. There was up to a 2.6-fold difference in the number of facts generated by different LLMs. Clinician review confirmed that while the generated facts were generally accurate, important details were often omitted. To enable reproducibility, we release FactEHR, a dataset encompassing 8,665 fact decompositions and 987,266 entailment pairs from four note types and three institutions, including 1,036 pairs annotated by seven clinical experts. The results highlight the need to improve LLM-based fact decomposition methods for clinical documents to support the use of LLMs for healthcare use.
6 Limitations
@@
First, our study focuses on four common note types, but many other specialized note types (e.g., ophthalmology) may yield even greater variability in fact decomposition, suggesting our error estimates are conservative. Second, we rely on a combination of open-source and proprietary models within a HIPAA-compliant environment, ensuring privacy but limiting reproducibility; we mitigate this by releasing our dataset. Thirdly, we did not control for prompt templates across models. Another limitation of this study is the lack of access to ground truth fact decompositions for clinical notes. This constraint forces us to rely on estimates and proxy measures, such as fact recall, to assess the completeness of model-generated fact decompositions. These metrics are inherently limited in their ability to capture the true completeness and accuracy of the decompositions. Lastly, since we examine only English documents, the generalizability of these findings to other languages remains uncertain.
Acknowledgements
@@
This work is generously supported by the Mark and Debra Leslie endowment for AI in Healthcare. We would also like to thank the Clinical Excellence Research Center (CERC) at Stanford for their support. AS thanks the Paul and Daisy Soros Foundation and Knight Hennessy Scholars for their support.
@@ A.1 Data Sources for Clinical Notes
For FactEHR, we sampled 2,168 clinical notes from three de-identified research datasets. We describe each of these datasets in detail:
@@ A.2 Entailment Data Sources
(Williams et al., 2018b) The Multi-Genre Natural Language Inference corpus is a crowd-sourced collection of 433k sentence pairs annotated with textual entailment information. The labels include entailment, neutral and contradiction.
@@ Appendix B Note Types
Appendix C Hyperparameters
@@
Appendix D Annotation Guidelines
@@
Appendix E Compute Environment
Experiments are performed in a HIPAA compliant APIs and local on-prem university compute environment using x Nvidia V100 GPUs. All compute environments supported HIPAA-compliant data protocols.
Appendix F Sentence Tokenization
@@
Appendix G Supplementary Results
G.1 Entailment model benchmarking
We report the performance of five LLMs (GPT-4o, GPT-4o-mini, Gemini-1.5-flash, Llama3-8b-Instruct, Llama3-70b-instruct) across five NLI datasets (FactEHR development set, MedNLI, SciTail, MultiNLI, and SNLI) in Table 6.
Dataset Model N Unparsable P R F1 Acc.
@@ G.1 Entailment model benchmarking
Table 6: Entailment evaluation benchmarking results. Note that Gemini-1.5 flash failed to process some examples due to content moderation ("N" column). Responses whose JSON output were not parsable were coerced to "0". Bolded are the highest numbers (F1 and accuracy) of a model for each dataset. Our results show that GPT-4o performs well on four out of five benchmarking datasets, including the FactEHR validation set. Therefore, we selected GPT-4o as the final LLM entailment judge.
G.2 Fact Atomicity

### v2

@@ 5.1 RQ1 — Fact Decomposition
We examine the length and similarity of fact decompositions across LLMs, as well as the atomicity of source sentences.
@@
Figure 3: Distribution of coefficient of variation across documents. Higher values indicate higher variance in number of facts across LLMs.
Figure 4: Sentence atomicity by note type and LLM, shown as the normalized distribution of the estimated number of facts-per-sentence
@@
Using (EMD) with a cosine cost function, we found that GPT-4o and o1-mini produced the most similar fact decompositions, closely followed by Gemini-1.5, whose pairwise EMD values against these models were consistently within 0.02 units of each other (Figure 5). In contrast, Llama3’s decompositions were notably more divergent, with pairwise EMD scores roughly 0.03–0.05 units higher. Compared to the original source sentences, GPT-4o decompositions had the lowest mean EMD, while Llama3’s were highest. Differences were also influenced by note type: procedure notes showed minimal variation due to their structured format, nursing notes had slightly higher EMDs reflecting more dynamic content, and discharge summaries—being lengthy and complex—exhibited the highest EMD values.
Sentence Atomicity
Overall, o1-mini produced the most facts per sentence, followed closely by GPT-4o, both maintaining similar rates of zero-fact sentences (Figure 4). Gemini-1.5 had more zero-fact sentences than GPT-4o and o1-mini, but fewer than Llama3.
Model Name Procedure Note Nursing Note Progress Note Discharge Summary
GPT-4o 1.45 $\pm$ 0.03 2.37 $\pm$ 0.14 1.28 $\pm$ 0.03 1.30 $\pm$ 0.02
o1-mini 1.45 $\pm$ 0.03 2.42 $\pm$ 0.13 1.46 $\pm$ 0.03 1.55 $\pm$ 0.03
Gemini-1.5 1.12 $\pm$ 0.03 2.05 $\pm$ 0.12 1.02 $\pm$ 0.02 0.98 $\pm$ 0.02
Llama3 8B 0.89 $\pm$ 0.03 1.58 $\pm$ 0.12 0.73 $\pm$ 0.04 0.60 $\pm$ 0.04
Table 2: Average facts per sentence by model and note type. Red indicates fewer facts than sentences, suggesting the LLM may fail to capture all facts.
Llama3 produced facts for only about 20% of sentences in progress notes and discharge summaries, and, except for nursing notes, generated fewer facts than sentences (Table 2).
Procedure Note Nursing Note
PP* RR* $R*_{w}$ F1F1* $F1*_{w}$ PP* RR* $R*_{w}$ F1F1* $F1*_{w}$
GPT-4o 98.5 78.7 92.4 86.5 95.0 97.5 88.9 92.2 92.3 94.2
o1-mini 97.8 78.4 93.0 86.2 95.4 96.6 88.5 93.3 91.6 94.4
Gemini-1.5 95.9 64.2 85.7 77.0 91.7 93.2 77.3 83.7 84.4 88.3
Llama3-8B 84.2 49.4 73.7 62.0 82.3 84.1 56.6 69.9 65.1 75.6
Progress Note Discharge Summary
PP* RR* $R*_{w}$ F1F1* $F1*_{w}$ PP* RR* $R*_{w}$ F1F1* $F1*_{w}$
GPT-4o 97.4 78.1 85.2 85.8 89.9 97.0 79.0 86.9 86.8 91.1
o1-mini 95.2 79.7 85.7 86.6 89.7 94.8 81.1 88.2 86.7 90.0
Gemini-1.5 96.0 66.2 75.2 77.9 83.9 96.5 65.4 76.3 76.9 83.8
Llama3-8B 83.4 41.9 57.4 53.8 67.0 84.6 38.9 57.3 50.4 67.2
[…]
In summary, quantifying this variability is relevant for two reasons: (1) Fact Verification Approaches: Current workflows use LLM-generated decompositions to break claims into sub-claims for verification. However, our analysis reveals inconsistencies across LLMs. Ideally, facts should be atomic, addressing one property or action at a time. When decompositions fail to achieve this, they introduce uncertainty and lead to varying conclusions, and (2) Clinical Document Types: Clinical documents vary in length, structure, and complexity, challenging LLM-based decomposition. Radiology reports are concise, while discharge summaries are lengthy and complex. Procedure notes require detailed step-by-step breakdowns. LLMs struggle with longer, intricate documents, often producing incomplete decompositions. Addressing these challenges is key to improving LLM performance in clinical documentation.
5.2 RQ2 — Entailment
GPT-4o achieve the best F1 scores on four of the five benchmark datasets, including the FactEHR development set in table 10 (Appendix I), and was therefore selected as the entailment judge for all 987,266 entailment pairs in the FactEHR dataset. From these, we randomly sampled 1,036 pairs for manual annotation by seven clinical experts. On doubly annotated entailment pairs, Fleiss’ Kappa was 0.67, indicating substantial inter-rater agreement. Under this evaluation, GPT-4o achieves a recall of 0.96 and precision of 0.86 (Appendix I, Table 11), closely mirroring its performance on the FactEHR development set. Overall, these findings suggest that GPT-4o serves as a reasonable proxy for human judgment in entailment assessments.
5.3 RQ3 — Fact Verification
Across all note types, fact precision remained consistently higher and less variable than fact recall—indicating that while the generated facts were generally accurate, many relevant facts were omitted, resulting in greater variability in recall and aligning with our qualitative findings (Appendix I, Figure 8).
Qualitative Expert Evaluation
Expert reviewers assessed overall completeness of fact decompositions and the correctness, independence, and atomicity of individual facts. Completeness, defined as capturing all information in the source, is low: only 5% (1/20) of decompositions meet this criterion. For individual facts, reviewers judge whether each fact accurately reflected the source (correctness), stood on its own (independence), and represented a minimal unit of information (atomicity), we present detailed annotation guidelines in Appendix F.
@@ 5.3 RQ3 — Fact Verification
Across models, Llama3 had the highest rate of incorrect facts. Discharge summaries yield the least number of atomic facts, reflecting their complexity, while progress and procedure notes contain more granular and independent facts. Detailed results for GPT-4o are shown in Figure 6; additional results for other models and a summary of key metrics are reported in Appendix J and Table 4, respectively.
@@
Overall, fact decompositions by Llama3 on long notes are subpar, specifically they are either incomplete or they get into infinite generation loops. Fact decompositions by GPT-4o are better, but also not complete when there are long notes. Fact decomposition appears to be easier for shorter notes than super long/dense notes. Even though some long notes get decomposed to 200+ facts, they are still missing content that was not decomposed. Llama3 and GPT-4o seem to have a greater tendency to produce incomplete statements which is measured as a non-independent statement as compared to Gemini and o1-mini.
6 Discussion
@@
In this study, we evaluate four LLMs’ ability to decompose over 2,000 real-world, EHR-derived clinical notes of four types into independent facts. We evaluate the fact decompositions with an LLM-as-a-Judge (validated by expert evaluation), as an entailment task. For calculating Fact-Precision, we check if the generated fact (the hypothesis) logically follows from original clinical note (the premise) and vice versa for Fact-Recall [^4] Detailed explanation in Section 3.5. To enable reproducibility and further research in this direction, we will release FactEHR, a dataset encompassing 8,665 fact decompositions and 987,266 entailment pairs from four note types and three institutions, including 1,036 pairs annotated by seven clinical experts. The dataset was specifically designed to address a gap in the literature regarding the accuracy of fact decomposition in clinical text. While fact decomposition is a critical first step for many fact-checking and verification methods, such as FactScore (Min et al., 2023), there is limited research on its effectiveness, particularly when applied to both human- and LLM-generated clinical text. Our work aims to fill this gap by evaluating LLM performance on this foundational task, which is often overlooked in existing studies. Additionally, we examine the length and similarity of the decompositions and find that there was up to a 2.6-fold difference in the number of facts generated by different LLMs. Clinician review confirmed that while the generated facts were generally accurate, although, important details were often omitted.
While choosing the LLMs for fact decomposition task, we focused on evaluating a representative set of high-performing general-domain models, as these have demonstrated broad utility and scalability for real-world applications. That said, we acknowledge the importance of benchmarking against additional domain-specific models and discuss the need to evaluate them in future work. We also emphasize that the methodology proposed in this study is model-agnostic and can be applied to evaluate the performance of other models, including ones that are medically pretrained. The release of the FactEHR dataset directly supports future evaluation of LLMs by the community.
In the future, we hypothesize that the dataset could be used to fine-tuning a smaller language model for fact-decomposition tasks, which could leverage the training and human-evaluated data splits from FactEHR to achieve higher quality decompositions tailored to specific needs. The results highlight the need to improve LLM-based fact decomposition methods for clinical documents to support the use of LLMs for healthcare use.
Limitations
@@
There are several limitations to this work. First, our study focuses on four common note types, but many other specialized note types (e.g., ophthalmology) may yield even greater variability in fact decomposition, suggesting our error estimates are conservative. Second, we rely on a combination of open-source and proprietary models within a HIPAA-compliant environment, ensuring privacy but limiting reproducibility; we mitigate this by releasing our dataset, prompts and code. Third, lack of access to ground truth fact decompositions for clinical notes. This constraint forces us to rely on estimates and proxy measures, such as Fact-Recall, to assess the completeness of model-generated fact decompositions. These metrics are inherently limited in their ability to capture the true completeness and accuracy of the decompositions. Lastly, since we examine only English documents, the generalizability of these findings to other languages remains uncertain. Additionally, there could be potential risks with the data. There could be data privacy concerns due to the de-identified but still potentially identifiable patient information, potential biases in the data. We acknowledge the potential for bias when using LLMs as judges. To mitigate this risk, we incorporated several safeguards; (a) LLM predictions are bench-marked against the human-annotated subset to ensure alignment and identify areas where LLMs might deviate (b) statistical power calculations guided the size of the human-labeled subset, ensuring it was large enough to rigorously validate LLM performance (c) our study explicitly discusses the limitations of LLM-based evaluation and highlights the importance of human-annotated benchmarks for future work.
In conclusion, we believe there is significant (and growing) utility to releasing large-scale synthetic and partially synthetic datasets in conjunction with LLM-based evaluators, reflected here by our FactEHRdataset and NLI clinician-curated dataset. This is especially important for academic labs that lack access to frontier model APIs that are HIPAA-compliant, creating barriers to exploring clinical text problems. Curating a sample of important clinical document types across multiple institutions and providing frontier LLM fact decompositions directly enables the research community to explore key problems in synthetic data quality evaluation, preference alignment, model distillation and more. By combining human-verified annotations with scalable LLM-based evaluations, our approach achieves a balance between scalability and the rigor needed for clinical applications, also keeping LLM bias in check (Li et al., 2024). This also ensures reliability as we rigorously validated the performance of LLMs by benchmarking them against a carefully selected subset of manually labeled examples. We also explicitly discusses the limitations of LLM-based evaluation earlier in this section and highlights the importance of human-annotated benchmarks for future work.
Acknowledgments
@@
This work is generously supported by the Mark and Debra Leslie endowment for AI in Healthcare. The authors also thank the Clinical Excellence Research Center (CERC) at Stanford for their support. AS thanks the Paul and Daisy Soros Foundation and Knight Hennessy Scholars for their support.
@@ A.1 Data Sources for Clinical Notes
For FactEHR, we sampled 2,168 clinical notes from three de-identified research datasets. We describe each of these datasets in detail and summarize the number of notes from each of them in table 5 and token statistics in table 6. We describe each source in detail below:
@@ A.2 Entailment Data Sources
(Williams et al., 2018) The Multi-Genre Natural Language Inference corpus is a crowd-sourced collection of 433k sentence pairs annotated with textual entailment information. The labels include entailment, neutral and contradiction.
@@ Appendix B Note Types
Appendix C Choice of Models
While we acknowledge the potential value of domain-specific models, numerous studies have shown that general-domain models often outperform medically pretrained(Jeong et al., 2024; Van Veen et al., 2024; Ceballos-Arroyo et al., 2024), or fine-tuned models across a range of tasks such classification, information extraction, and summarization, including our initial experiments.
Note Type MedLM Gemini-1.5
Discharge summary 58.5 64
Nursing note 77.4 79.8
Progress note 74.8 77.5
Table 7: Comparison of scores of MedLM and Gemini. We report Fact-F1 scores. Gemini-1.5 consistently outperformed MedLM in our evaluations.
Appendix D Hyperparameters
@@
Appendix E Entailment Annotation Guidelines
@@
Appendix F Qualitative Annotation Guidelines
Expert reviewers assessed overall completeness of fact decompositions and the correctness, independence, and atomicity of individual facts.
Completeness: The set of fact decompositions are defined complete if the facts capture all information present in the source document.
  1. 1.
    : 1: If all of the information in the note is covered in the decomposed fact list
  2. 2.
    : 0 if any part of the note is missed by decomposed facts.
  3. 3.
    Example: Note: "No evidence of pneumothorax.
    Left pleural effusion increased from prior scan."
    Facts: "No pneumothorax. There is a pleural effusion."
    Missing: "Pleural effusion is increased from last scan."
    Score: 0
For individual facts, reviewers judge on three criterion:
Atomicity: A fact is atomic if it represented a minimal unit of information which cannot be decomposed further.
  1. 1.
    1: The sentence presents exactly one fact, statement, or concept without unnecessary complexity.
  2. 2.
    0: The statement includes multiple facts, making it difficult to evaluate or extract key information.
  3. 3.
    Example: Note: "No evidence of pneumothorax.
    Left pleural effusion increased from prior scan."
    Facts: "Pleural effusion is left sided and increased from last scan."
    Score: 0
Correctness: A fact is correct if it accurately reflects the contents of the source.
  1. 1.
    1: The fact presented accurately reflects the information in the note.
  2. 2.
    0: The statement includes multiple facts, making it difficult to evaluate or extract key information.
  3. 3.
    Example: Note: "No evidence of pneumothorax.
    Left pleural effusion increased from prior scan."
    Facts: "There is a large pneumothorax."
    Score: 0
Independence: A fact is independent if it can be understood on its own.
  1. 1.
    1: The fact presented accurately reflects the information in the note.
  2. 2.
    0: The statement includes multiple facts, making it difficult to evaluate or extract key information.
  3. 3.
    Example: Note: "No evidence of pneumothorax.
    Left pleural effusion increased from prior scan."
    Facts: "It is left sided."
    Score: 0
Appendix G Compute Environment
Experiments are performed in a HIPAA compliant APIs and local on-prem university compute environment using 4 Nvidia H100 GPUs. All compute environments supported HIPAA-compliant data protocols.
Appendix H Sentence Tokenization
@@
Note Type Junk Partial Examples
Procedures 9.0% 6.0% ___, _________, I., *, (Over), $\mid$, _______
Appendix I Supplementary Results
I.1 Validating LLM-as-a-Judge
A key objective of this work was to evaluate the “LLM-as-a-judge" component within a fact decomposition verification framework. Despite the critical role of this initial step in fact verification pipelines, it has not been rigorously benchmarked in clinical text or across diverse clinical document types. The scale of human-annotated labels used in this study was driven by the need to validate our proposed NLI-based LLM-as-a-judge. To determine an appropriate sample size for these human-provided labels, we performed power calculations. These calculations ensured that our manually labeled dataset was sufficiently powered to accurately estimate sensitivity, specificity, and other performance metrics, using parameters derived from NLI benchmarking results on the FactEHR development set.
With these parameters, annotating 250 entailment pairs per note type was sufficient to provide high-confidence estimates (99% confidence level overall and  80% confidence level per note type). This resulted in a total of 1000 uniquely annotated entailment pairs across the four note types. Additionally, we included 200 duplicate annotations (50 per note type) to assess inter-annotator agreement, ensuring we could reliably estimate kappa with high confidence. This brought the total for this annotation tranche to 1200 entailment pairs, including both unique and duplicate annotations. Combined with the 2468 entailment pairs already annotated, our dataset now comprises approximately 3500 human-annotated entailment pairs, representing one of the largest datasets of its kind in the clinical domain.
Metric Performance
Expected sensitivity 95%
Expected specificity 87%
Prevalence of entailed claims 70%
Table 9: Performance of LLM-as-a-Judge model on the 250 entailment pairs per note type, they provide high-confidence estimates.
Dataset Model N Unparsable P R* F1 Acc.
@@ I.1 Validating LLM-as-a-Judge
Table 10: Entailment evaluation benchmarking results. Note that Gemini-1.5 flash failed to process some examples due to content moderation (“N" column). Responses whose JSON output were not parsable were coerced to “0". Bolded are the highest numbers (F1 and accuracy) of a model for each dataset. Our results show that GPT-4o performs well on four out of five benchmarking datasets, including the FactEHR validation set. Therefore, we selected GPT-4o as the final LLM entailment judge.
I.2 Fact Atomicity